OpenStudy (automaticloveletter):

Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 5/3 , 4/3 -5/3 , -4/3 -3/5 , -3/4 3/4 , 3/5

OpenStudy (anonymous):

x + y = - 3 move y to the other side x+ y - y = - 3 - y ^^^^ those cancel out x = - 3 - y now input "x = - 3 - y" into "y = 2x + 2" and tell me what you get :)

OpenStudy (anonymous):

if you're having trouble doing that, just tell me and I can help you step by step

OpenStudy (automaticloveletter):

2*(-3)+2=(-4)?

TheSmartOne (thesmartone):

Why are you doing it the hard way...

TheSmartOne (thesmartone):

x + y = −3 y = 2x + 2 Substitute \(\sf y=2x+2\) into \(\sf x+y=-3\)

TheSmartOne (thesmartone):

So: \(\sf\color{red}{y}=2x+2\) \(\sf x+\color{red}{y}=-3\)

TheSmartOne (thesmartone):

Substitute it, you get: \(\sf x+2x+2 = -3\)

OpenStudy (anonymous):

That's just the way that I am used to doing it. I didn't mean to offend you. Everyone has their own way of solving things differently that works for them and that leads to the same answer.

OpenStudy (automaticloveletter):

-4/3 right

TheSmartOne (thesmartone):

that's wrong...

OpenStudy (automaticloveletter):

Well that's odd, considering I got full points on my question.